OpenStudy (anonymous):

2x=x2-5x -- ----- 5 5x 2x/5=xsquare-5x/5x

OpenStudy (mridul):

10x^2 = 5x^2- 25x 5x^2 = -25x x^2 = -5x x^2 + 5x = 0 (x + 5)x = 0 x+5=0 x=-5

OpenStudy (anonymous):